OpenStudy (faman39):

Russell deposited $650 in a savings account that earns 3% simple annual interest. If he makes no other deposits or withdraws, what will the balance of Russell's account be after 6 years?

OpenStudy (anonymous):

You are going to fail your exams if you don't try to figure any of this stuff out.

OpenStudy (anonymous):

The answer is E

OpenStudy (bahrom7893):

What's a simple annual interest?

OpenStudy (faman39):

Can u do another?

OpenStudy (bahrom7893):

theyre all simple annual interest. hold on let me look up the formula.

OpenStudy (bahrom7893):

I = P*r*t I = interest P = Principal t = time r = rate

OpenStudy (bahrom7893):

Russell deposited $650 in a savings account that earns 3% simple annual interest. If he makes no other deposits or withdraws, what will the balance of Russell's account be after 6 years? Hm what are principal and other stuff in this equation?

OpenStudy (anonymous):

find interest P* R* T P=650 R=.03 T=6 and add it to principle of $650

OpenStudy (bahrom7893):

imran to the rescue lol, sorry faman i haven't really done applied math. Like if u give me the formulas and what stands for what, I'll get u the answer. Otherwise, I'm pretty much helpless lol